OpenStudy (lacedupcountry):

Consumer math problem: Can someone help me, I can't figure out what I'm doing wrong. I'll post the questions below! (Will Give medals!)

OpenStudy (lacedupcountry):

You worked from 8:30 am to 2:15 pm with a 1/2 hour lunch break. How long did you work? A.5.25 B.5.75 C.6.25 D.6.75 I added all thw hours up and keep getting 5.45 hours, i don't know what i'm doing wrong! Also lunch breaks arent counted towards your pay if that helps

OpenStudy (dumbcow):

8:30 to 2:15 is 5 hrs 45 min subtract 30 min lunch ---> 5 hr 15 min = 5.25 hrs 15/60 = 1/4 = .25

OpenStudy (lacedupcountry):

Oh okay i see what i did wrong here thanks!

OpenStudy (dumbcow):

yw :)